Startup of DX AdSec and Alignment of 4D 20170823

From 9am to 4pm MST
  • Took DX AdSec cover off
  • Powered on 110V in AO Cabinet
  • 4D started.
    • Tertiary set to values determined in June
      • -1500, -600, 0, 39.6 (Seems like large Tip/Tilt)
    • Pencil beam found on center of AdSec.
    • Installed Divergent lens
  • Powered on DX AdSec (using aoroot AO Supervisor)
    • Found 42 um bump at 4:00, Act 596
    • TSS not enabled
  • Moved to Zenith
    • Still 42 um
    • Set Shell, no problems *Rested Shell
      • 42 um bump went away
      • 32 um bump at 12:00, act 39
      • Set and Rested several times, still 32 um bump at 12:00
      • Installed act_add AO Supervisor
        • Found that act_add still had ff_matrix.fits set to NULL
        • Set back to FF matrix file
        • Set Shell. Jumping actuator 559. Maybe will remove from use.
    • 4D spot found on back of RR (near pinhole) looking from level 10
      • Current position of hexapod (set in early July when Hexapod turned off)
        • 0,0,0,0,-10,0
      • Out of focus
      • Will try to set hexapod to 0,0,0,0,0,0 once it's powered on and homed
      • Will probably not try to find the spot at Horizon, but move back to Zenith and look from level 10
  • Move back to Horizon
    • Remove TSS disable power supply
    • Test TSS will engage when anemometer process is powered off

5:48 Search for spot on back of RR
  • hexapod value 0,0,0,0,0,0
  • Phil can see a spot near the pinhole
    • may be some light entering pinhole
  • Move out of focus to better see spot on pinhole
    • 0,0,0,0,5,0
  • Found Fringes. Very Comatic
    • 0,0,-400,-460, 5, 0

7:38 Tertiary Position
  • 1500, -600, 0, 39.60

7:47 Secondary Hexapod
  • -3.6, -4.03, 5.15, 0, -806, 0
  • -4.7, -4.4, 5.15, 40, -916, 0

9:30 Apply astigmatism to AdSec
  • Z5 = -1.0
  • Z6 = -1.7

10:05 4D Connection
  • vncviewer 192.168.18.180

10:56 Remove actuator 559
  • Jumps to high current probably because of jump in position
  • current file is bad_actuator.list.2017-06-30T08:32:00
  • new file is backup/bad_actuator.list.2017-08-24T06:03:28
    • Also removed from pos do-not-use list actuators
      • 340, 453, 484, 452
      • descriptions just said high forces resulting in poor flats
      • test if this is still a problem
    • Added more actuators to use
      • 557

12:08
  • Hexapod -4.915, -4.375, 5.15, -35, -939
  • On adsec
    • Z2 = 7.0
    • Z3 = 0.0
    • Z5 = -0.6
    • Z6 = -1.2

-- %USERSIG{DougMiller - 2017-08-23}%
Topic revision: r5 - 24 Aug 2017, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback